OpenStudy (anonymous):

Solve for X : log[6](x+4)-log[6](x-4)=log[6]^(3)

OpenStudy (anonymous):

\[\log _{6}(x+4)-\log _{6}(x-4)=\log _{6}^{3}\]

OpenStudy (anonymous):

already tried that. it didnt give me a solution

OpenStudy (angela210793):

is it this?